# Adaptive Layer-skipping in Pre-trained LLMs

### v1

@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in various tasks, including translation [44], code generation [16], and math reasoning [6]. Currently, LLMs typically generate each token by performing a full forward pass through all Transformer decoder layers. However, such a uniform allocation is counter-intuitive, as simpler tasks intuitively require fewer computational resources, whereas more complex tasks demand more extensive processing. This uniform allocation not only results in computational inefficiencies but may also contribute to overfitting in LLMs. To improve the efficiency of token generation, various layer-skipping techniques have been proposed. One line of research leverages statistical information, such as the difference between layer inputs and outputs, to identify and skip less important layers [26, 13, 43]. Another collection of methods involves early-exit [37, 18, 34], where a confidence measure or a router dynamically determines whether to bypass all subsequent layers at an intermediate point of LLM. While these techniques have successfully reduced computational costs, they overlook a fundamental question: How do computational demands vary across the generation of different tokens
@@
Figure 1: Token depth map illustrating layer-skipping patterns when applying FlexiDepth to Llama-3-8B-Instruct. The light-to-dark blue gradient indicates layer usage ranging from 16 to 32.
@@
Beyond performance gain, FlexiDepth can reveal layer allocation patterns of pre-trained LLMs, addressing the question of how computational demands vary when generating different types of tokens. To demonstrate this phenomenon, we created a colored map that shows the number of layers used to generate each token, and call it "token depth map". This token depth map visualization, as shown in Figure 1 (left), reveals that for different language generation tasks, summarization typically utilizes more layers on average than extractive question answering. Similarly, in mathematical reasoning tasks like addition (Figure 1, right), tokens on the left-hand side of equations require fewer layers to generate compared to those on the right-hand side. This type of adaptive allocation appears to resonate with human intuition. For instance, in language tasks, humans typically exert more effort in summarization to condense the entire context into a concise form. In contrast, extractive question answering often relies on retrieval and extraction, making it easier to identify the answer. Likewise, in mathematical tasks, deriving the right-hand side of an equation involves computation that consumes more steps. To inspire further research along this direction, we compile a dataset documenting the layer allocation behaviors of FlexiDepth across various tasks.
@@ 2.2 Attention skipping
In the skipping path of a FlexiDepth layer (Figure 2 right), the input hidden states simply bypass the attention module via a shortcut. WhileHowever, this naive approach reduces computational overhead, it risks losing critical contextual informleads to a significant loss of contextual information, resulting in performance degradation. The attention mechanism depends onleverages query, key, and value vectors to enable tokens to attend to one another,each other, thereby capturing long-range dependencies essential for generation. When aFor hidden states that skips the attention module entirely, its, their corresponding key and value vectors are.. not generated generated, preventing subsequent tokens from accessing its informationttending to them. We show an example in Figure 3 to describe this.
As illustrated in Figure 3 (left), in the vanilla attention mask, every token can attend to all preceding tokens. For instance, query vector $q_{4}$ attends to keys $k_{0}$ through $k_{4}$. However, when employing FlexiDepth without maintaining the key-value cache (KV-cache), skipping hidden states such as $x_{1}$ and $x_{3}$ results in missing keys $k_{1}$ and $k_{3}$, as depicted in the attention mask labeled ’No KV Cache’ (Figure 3, middle). Consequently, future tokens like $q_{4}$ lose access to the contextual information stored in these hidden states.
To address this issue, we propose a simple solution: we still compute the KV-cache for hidden states that skip the attention module. Specifically, we still bypass the computation of query vectors (e.g., $q_{1}$ and $q_{3}$) and their associated scaled dot-product operations for the skipped states $x_{1}$ and $x_{3}$. However, we retain the calculation of their corresponding keys $k_{1},k_{3}$ and values $v_{1},v_{3}$, as demonstrated in the attention mask labeled ’KV Cache’ (Figure 3, right). This ensures that future generation maintains access to the entire context, preserving the integrity of autoregressive generation.
@@ 3.2 Main results
3.3 FlexiDepth on different language models
We evaluate FlexiDepth on language models of varying sizes. These models are trained under the identical configuration and dataset described in our implementation details. All of the models are instruction-tuned models. Larger models, specifically Llama-2-13B-Instruct and Llama-3-8B-Instruct, exhibit greater number of layers skipped compared to the smaller Qwen-2.5-3B-Instruct [38]. Concretely, the Llama-2-13B-Instruct and Llama-3-8B-Instruct skip approximately 7 and 6 layers respectively across various tasks, whereas the Qwen-2.5-3B-Instruct skips only about 1 to 2 layers. This pattern suggests that larger models inherently possess higher redundancy, thus allowing more aggressive layer skipping without significant performance deterioration. These observations highlight the substantial potential of adaptive layer skipping methods when applied to larger language models.
@@ 6 Conclusion
We present FlexiDepth, an approach to dynamically skip layers in pre-trained large language models. It achieves state-of-the-art performance without altering the original model parameters. FlexiDepth offers valuable insights into the varying computational demands of token generation. We constructed a dataset capturing these allocation patterns to encourage further study of computational needs of tokens, thereby advancing the development of more adaptive and resource-efficient language models.

### v2

@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in various tasks, including translation [45], code generation [17], and math reasoning [6]. Currently, LLMs typically generate each token by performing a full forward pass through all Transformer decoder layers. However, such uniform allocation is counter-intuitive, as simpler tasks intuitively require fewer computational resources, whereas more complex tasks demand more extensive processing. This uniform allocation not only results in computational inefficiencies but may also contribute to overfitting in LLMs. To mitigate the inefficiencies of uniform allocation, various layer-skipping techniques have been proposed for acceleration. One line of research leverages statistical information, such as the difference between layer inputs and outputs, to identify and skip less important layers [27, 14, 44]. Another collection of methods involves early-exit [38, 19, 35], where a confidence measure or a router dynamically determines whether to bypass all subsequent layers at an intermediate point of LLM. While these techniques have successfully reduced computational costs, they overlook a fundamental question: How do computational demands vary across the generation of different tokens?
@@
Figure 1: DepthMap illustrating layer-skipping patterns when applying FlexiDepth to Llama-3-8B-Instruct. The light-to-dark blue gradient indicates layer usage ranging from 16 to 32.
@@
Beyond performance gain, FlexiDepth can reveal layer allocation patterns of pre-trained LLMs, addressing the question of how computational demands vary when generating different types of tokens. To demonstrate this phenomenon, we created a colored map that shows the number of layers used to generate each token, and call it "DepthMap". Figure 1 (left) shows that summarization typically requires more layers on average than extractive question answering. Similarly, Figure 1 (right) reveals that in mathematical reasoning tasks like addition, tokens on the left-hand side of equations need fewer layers to generate than those on the right-hand side. This type of adaptive allocation appears to resonate with human intuition. For instance, in language tasks, humans typically exert more effort in summarization to condense the entire context into a concise form. In contrast, extractive question answering often relies on retrieval and extraction, making it easier to identify the answer. Likewise, in mathematical tasks, deriving the right-hand side of an equation involves computation that consumes more steps. Shifting from token-specific patterns to broader layer usage, we observe that layer utilization follows a long-tail distribution, with early and final layers used more frequently than middle layers. To inspire further research along this direction, we compile a dataset documenting the layer allocation behaviors of FlexiDepth across various tasks.
@@ 2.2 Attention skipping
In the skipping path of FlexiDepth (Figure 2, right), the input hidden states simply bypass the attention module via a shortcut. However, this straightforward approach can lead to a significant loss of contextual information, resulting in performance degradation. The attention mechanism relies on query, key, and value vectors to enable tokens to attend to each other, capturing long-range dependencies for generation. When hidden states skip the attention module, their corresponding key and value vectors are not generated, preventing subsequent tokens from attending to them. This issue is illustrated in Figure 3 (middle), where in this example, hidden states $x_{1}$ and $x_{3}$ are routed to skip the layer. Naively skipping the attention module for these tokens results in missing key-value pairs (i.e., $k_{1},v_{1}$ and $k_{3},v_{3}$). Therefore, the subsequent tokens, such as $q_{4}$, can no longer attend to them.
To address this issue, we propose a straightforward solution: we continue computing the keys and values for hidden states that skip the layer. Specifically, while we omit the query vectors (e.g., $q_{1},q_{3}$) and their associated scaled dot-product operations, we retain the calculations of the corresponding keys ($k_{1},k_{3}$) and values ($v_{1},v_{3}$). This approach, as demonstrated in Figure 3 (right), ensures that future tokens maintain access to the full set of contextual information, preserving the integrity of autoregressive generation.
@@ 3.2 Main results
3.3 FlexiDepth on Different Language Models
We evaluate FlexiDepth across language models of varying sizes to assess its adaptability. All models, which are instruction-tuned, were trained using the same configuration and dataset as outlined in our implementation details. Larger models, namely Llama-2-13B-Instruct and Llama-3-8B-Instruct, skip more layers compared to the smaller Qwen-2.5-3B-Instruct [39]. Specifically, Llama-2-13B-Instruct and Llama-3-8B-Instruct skip an average of 6.42 and 6.25 layers, respectively, across benchmarks, while Qwen-2.5-3B-Instruct skips only about 1.49 layers. This trend indicates that larger models possess greater inherent redundancy, enabling more aggressive layer skipping without notable performance degradation. Given this trend, FlexiDepth can potentially be applied to even larger models to further exploit their redundancy. For instance, it could be applied to large mixture-of-experts models like DeepSeek-V3 [7], where experts are distributed across multiple servers during inference. By skipping entire layers, FlexiDepth could reduce cross-server communication overhead, enhancing inference efficiency of MoE models.
@@ 3.4 Layer allocation dataset
Figure 4: Percentage of tokens processed by transformer layers 17 to 32. The x-axis represents the layer index, and the y-axis represents the percentage of tokens processed by the layer.
3.5 Layer utilization pattern
To complement the token-centric analysis in prior subsections, we examine the layer utilization of FlexiDepth on Llama-3-8B-Instruct. We measure the percentage of tokens processed by each transformer layer across tasks outlined earlier: text generation (copying, summarization, continuation) and math reasoning (repeating, addition, product). Figure 4 visualizes the results, with the left plot showing language tasks and the right plot showing math reasoning. Layer utilization displays a bowl-like pattern, with higher token processing at the initial and final layers, and reduced usage in the middle layers. For language tasks, copying exhibits the deepest dip in middle-layer usage, while continuation involves the most middle layers, reflecting varying task complexity. Similarly, in math reasoning, repeating shows the least middle-layer usage, whereas product requires the most, consistent with increasing task difficulty. The overall layer usage distribution follows a long-tail pattern, where most tokens are processed by a few layers, and others are selectively skipped. We hypothesize that early layers primarily parse and understand input context, middle layers perform task-specific processing that varies with complexity, and final layers focus on decoding for coherent output generation. This functional specialization explains the bowl-like pattern: early and final layers are essential for nearly all tokens, while middle layers are skipped more frequently for simpler tasks that require less intermediate processing. While this explanation remains a hypothesis, future studies are needed to verify the specialized roles of different layers.
Model MMLU Hellaswag Winogrande GSM8K HumanEval CoQA Retain %
Llama-3-8B 0.673 0.706 0.744 0.679 0.299 0.784 100.0%
FlexiExit 0.676 0.741 0.756 0.681 0.220 0.787 96.9%
[…]
Table 3: Performance and skipping patterns of FlexiExit
3.6 Early exit
We extend FlexiDepth’s layer-skipping mechanism into an early-exit framework called FlexiExit, which allows tokens to exit at intermediate layers. In FlexiExit, we adapt FlexiDepth’s router and adaptor modules so that a token flagged for skipping at a given layer bypasses the attention modules in all subsequent layers and is processed solely by lightweight adaptors. We applied FlexiExit to the Llama 3 8B Instruct model, training it under the same conditions as FlexiDepth but with a reduced layer-skipping weight of $\alpha=3\times 10^{-4}$. As shown in Table 3, FlexiExit retains the performance of the original Llama-3-8B-Instruct, while skipping an average of 1.61 to 4.68 layers. The current design of the router and adapter remains effective in supporting early exit scenarios.
@@ 6 Conclusion
We present FlexiDepth, an approach to dynamically skip layers in pre-trained large language models. It achieves state-of-the-art performance without altering the original model parameters. FlexiDepth offers valuable insights into the varying computational demands of token generation. We constructed a dataset capturing these allocation patterns to encourage further study of computational needs of tokens, thereby advancing the development of more adaptive and resource-efficient language models. Future works can be done for the application of FlexiDepth in practice..
7 Acknowledgements
This work was partially supported by the BioPACIFIC Materials Innovation Platform of the National Science Foundation under Award No. DMR-1933487.